### v1

# Pushing Toward the Simplex Vertices: A Simple Remedy for Code Collapse in Smoothed Vector Quantization

Abstract
Vector quantization, which discretizes a continuous vector space into a finite set of representative vectors (a codebook), has been widely adopted in modern machine learning. Despite its effectiveness, vector quantization poses a fundamental challenge: the non-differentiable quantization step blocks gradient backpropagation. Smoothed vector quantization addresses this issue by relaxing the hard assignment of a codebook vector into a weighted combination of codebook entries, represented as the matrix product of a simplex vector and the codebook. Effective smoothing requires two properties: (1) smoothed quantizers should remain close to a onehot vector, ensuring tight approximation, and (2) all codebook entries should be utilized, preventing code collapse. Existing methods typically address these desiderata separately. By contrast, the present study introduces a simple and intuitive regularization that promotes both simultaneously by minimizing the distance between each simplex vertex and its $K$-nearest smoothed quantizers. Experiments on representative benchmarks—including discrete image autoencoding and contrastive speech representation learning—demonstrate that the proposed method achieves more reliable codebook utilization and improves performance compared to prior approaches.
1 Introduction
Vector quantization is a method for discretizing a continuous vector space (Gray,, 1984; van den Oord et al.,, 2017). It maps each vector in the continuous space to the nearest element of a finite set of representative vectors (a.k.a. a codebook). The resulting discrete representations are easier to manipulate and interpret than the original continuous forms, and have proven effective across diverse applications, including image generation (Esser et al.,, 2021; Ramesh et al.,, 2021; Rombach et al.,, 2022; Yu et al.,, 2022), speech recognition (Baevski et al., 2020a, ; Baevski et al., 2020b, ), and music generation (Hadjeres and Crestel,, 2020; Dhariwal et al.,, 2020).
When integrated into deep neural networks, however, vector quantization introduces a fundamental challenge: quantization is a non-differentiable operation that blocks gradient backpropagation (van den Oord et al.,, 2017). Accordingly, some approximation is required to enable learning through quantization.
One effective workaround is to smooth vector quantization (Jang et al.,, 2017). The selection of a codebook vector can be expressed as multiplying the codebook matrix by its corresponding onehot vector $(0,\dots,1,\dots,0)^{\mathsf{T}}$, whose nonzero entry indexes the chosen vector. Smoothed vector quantization relaxes this onehot vector to lie within the simplex $\Delta^{M-1}:=\{(p_{1},\dots,p_{M})\mid\sum_{m=1}^{M}p_{m}=1\}$, where $M$ denotes the number of codebook vectors. Consequently, differentiable mappings (e.g., softmax) become available and can be incorporated into neural networks.
[…]
Figure 1: (1) Four different distributions on the simplex $\Delta^{3-1}$. For effective smoothed vector quantization, samples should be concentrated near the vertices of the simplex (i.e., onehot-like vectors; orange), rather than centered (dark gray) or uniformly spread across the simplex (light gray). At the same time, each vertex must be neighbored by some samples to avoid code collapse (blue). (1) Maximizing the perplexity of the sample mean (Baevski et al., 2020b, ) penalizes code collapse but cannot discriminate among the other three distributions. (1) The proposed $K$-nearest neighbor (KNN) distance minimization ($K=8$) favors the desired vertex-concentrated distribution while also preventing code collapse.
To successfully approximate onehot quantizers, smoothed quantizers must be distributed around the vertices of the simplex (orange “o” in Figure 1). At the same time, they should not concentrate around a few vertices, leaving other codebook entries unused (blue “x”). This latter issue—called code collapse—has been identified as a major challenge in vector quantization (Dieleman et al.,, 2018; Baevski et al., 2020b, ; Dhariwal et al.,, 2020; Fifty et al.,, 2025). A widely adopted workaround for code collapse (in smoothed quantization) is to introduce an auxiliary learning objective that maximizes the entropy or perplexity of the mean of the smoothed quantizers (Dieleman et al.,, 2018; Baevski et al., 2020b, , see §2.3 for the formal definition). However, maximizing the entropy/perplexity of the mean can be achieved by various distributions, not only the desired vertex-neighboring ones. For example, both uniform and center-concentrated distributions have their mean at the simplex center (Figure 1), which also maximizes this objective (Figure 1). Accordingly, an additional mechanism is needed to tighten the smoothing (e.g., by adjusting the temperature parameter of the (Gumbel-)softmax; see §2.2).
Beyond this standard approach, however, there exists a simple and unified strategy for simultaneously tightening smoothed quantization and maximizing codebook usage: Why don’t we directly encourage clustering around all simplex vertices? Specifically, minimizing the distance between each simplex vertex and its $K$-nearest neighbors (KNNs) satisfies both desiderata at once (Figure 1). The present study investigates this intuitive yet underexplored approach, comparing it against existing alternatives on representative benchmarks. The results indicate that the proposed method enables the exhaustive usage of the entire codebook, even when other approaches suffer from code collapse.
The contributions of this work are summarized as follows:
  • Neural vector quantization is reformulated as a smoothing problem of onehot vectors. This simple reformulation has been absent in the literature, which traditionally framed vector quantization as an extension of variational autoencoding (Kingma and Welling,, 2014; Jang et al.,, 2017; van den Oord et al.,, 2017).
  • Under this reformulation, an effective regularization loss function is proposed to promote both tight smoothing and exhaustive code utilization. The method demonstrates robustness across different learning settings.
The remainder of this paper is organized as follows. §2 reviews related studies on vector quantization. §3 introduces the proposed method, which is then evaluated on representative benchmarks in §4. Finally, §5 discusses the results and the limitations of the proposed method.
2 Related Studies
The central challenge of vector quantization lies in its non-differentiability, which disrupts the backpropagation of gradients in neural networks. To address this, either the gradient computation (backward path) or the quantization itself (forward path) must be approximated.
2.1 Approximation in the Backward Path (Gradient Estimation)
One line of work retains the original non-differentiable quantization in the forward path but replaces the gradient computation in the backward path. Let $\mathbf{z}\in\mathbf{R}^{D}$ denote a pre-quantized feature vector, and $\{\mathbf{q}_{1},\dots,\mathbf{q}_{M}\}\subset\mathbf{R}^{D}$ the set of quantized vectors. Quantization maps $\mathbf{z}$ to its “closest” codebook entry according under a distance metric $\mathcal{D}$: i.e., $\mathbf{z}\mapsto\mathbf{q}_{\iota(\mathbf{z})}$ where $\iota(\mathbf{z}):=\operatorname{argmin}_{m}\mathcal{D}(\mathbf{z},\mathbf{q}_{m})$. Consequently, the partial derivatives $\frac{\partial q_{\iota(\cdot),i}}{\partial z_{j}}$ are ill-defined and must be approximated.
A canonical approximation—known as the straight-through estimation (STE)—replaces the ill-defined Jacobian with the identity matrix (van den Oord et al.,, 2017):
$\frac{\partial q_{\iota(\cdot),i}}{\partial z_{j}}\approx\begin{cases}1&i=j\\ 0&\text{otherwise}\end{cases}$ (1)
Algorithmically, the STE is implemented using the $\mathtt{detach}$ operation, which excludes its argument from gradient computation:
$\operatorname{STE}(\mathbf{q}_{\iota(\mathbf{z})},\mathbf{z})$ $:=\mathbf{q}_{\iota(\mathbf{z})}+\mathbf{z}-\mathtt{detach}(\mathbf{z})$ (2)
Here, $z-\mathtt{detach}(z)$ evaluates to zero, while the gradient with respect to $z$ can still be propagated through the first term.
More recently, Fifty et al., (2025) proposed an alternative gradient approximation, and demonstrated its empirical superiority over the STE:
$\textsc{RE}(\mathbf{q}_{\iota(\mathbf{z})},\mathbf{z})$ $:=\mathtt{detach}\left(\frac{\|\mathbf{q}_{\iota(\mathbf{z})}\|}{\|\mathbf{z}\|}R\right)\mathbf{z}$ (3)
where $R$ is the rotation matrix aligning $\mathbf{z}$ to $\mathbf{q}_{\iota(\mathbf{z})}$, [^1] The rotation matrix is given by $R=I-2\hat{\mathbf{r}}\hat{\mathbf{r}}^{\mathsf{T}}+2\hat{\mathbf{q}}_{\iota(\mathbf{z})}\hat{\mathbf{z}}^{\mathsf{T}}$, where $\hat{\mathbf{v}}:=\mathbf{v}/\|\mathbf{v}\|$ is the L2-normalization of vector $\mathbf{v}$, and $\mathbf{r}:=\hat{\mathbf{q}}_{\iota(\mathbf{z})}+\hat{\mathbf{z}}$. and $\frac{\|\mathbf{q}_{\iota(\mathbf{z})}\|}{\|\mathbf{z}\|}$ rescales the rotated vector to match the amplitude of $\mathbf{q}_{\iota(\mathbf{z})}$. In this formulation, the Jacobian of the quantization is approximated by the scaled rotation matrix:
$\frac{\partial q_{\iota(\cdot),i}}{\partial z_{j}}\approx\frac{\|\mathbf{q}_{\iota(\mathbf{z})}\|}{\|\mathbf{z}\|}R$ (4)
2.2 Approximation in the Forward Path (Smoothing)
An alternative approach approximates the forward quantization itself. Using the onehot representation $\mathbf{e}_{m}$ of the code index $m$, quantization can be expressed as:
$\mathbf{q}_{\iota(\mathbf{z})}$ $=Q\mathbf{e}_{\iota(\mathbf{z})}$ (5)
where $Q:=(\mathbf{q}_{1},\dots,\mathbf{q}_{M})$. Smoothed quantization extends the possible range of $\mathbf{e}_{m}$ to the simplex $\Delta^{M-1}$. As noted in §1, effective learning requires smoothed quantizers $\mathbf{p}\in\Delta^{M-1}$ to concentrate near the vertices of the simplex (i.e., $\mathbf{p}\approx\mathbf{e}_{m}$ for some $m$).
A widely studied instance of smoothed quantization is Gumbel-softmax sampling (Jang et al.,, 2017). Given assignment probabilities $\pi_{m}$ of $\mathbf{z}$ to the $m$-th code—typically log-proportional to their dot-product $\mathbf{q}_{m}^{\mathsf{T}}\mathbf{z}$—categorical sampling can be implemented using Gumbel samples $g_{m}=-\log(-\log u_{m})$ with $u_{m}\sim\operatorname{Uniform}(0,1)$:
$\iota(\mathbf{z})\sim\operatorname{Categorical}(\pi_{1},\dots,\pi_{M})$
$\Leftrightarrow$ $\iota(\mathbf{z})=\underset{m}{\operatorname{argmax}}(g_{m}+\log\pi_{m})$ (6)
Replacing $\operatorname{argmax}$ above with $\operatorname{softmax}$ yields a smoothed quantization:
$p_{m}=$ $\frac{\exp\left(\left(g_{m}+\log\pi_{m}\right)/\tau\right)}{\sum_{m^{\prime}}\exp\left(\left(g_{m^{\prime}}+\log\pi_{m^{\prime}}\right)/\tau\right)}$ (7)
where lowering the temperature parameter $\tau$ produces a tighter approximation of categorical sampling.
The Gumbel-softmax sampling can also be combined with hard quantization using the STE:
$\operatorname{STE}(\mathbf{e}_{\iota(\mathbf{z})},\mathbf{p})=\mathbf{e}_{\iota(\mathbf{z})}+\mathbf{p}-\mathtt{detach}(\mathbf{p})$ (8)
2.3 Regularization
Beyond approximation strategies, vector quantization also requires auxiliary regularization losses to ensure effective training. For example, the STE alone does not guarantee alignment between pre-quantized features and codebook entries. This alignment is instead fostered by the following regularization loss, $\mathcal{L}_{\text{reg}}$ (van den Oord et al.,, 2017; Fifty et al.,, 2025):
$\mathcal{L}_{\text{total}}$ $=\mathcal{L}_{\text{main}}+\mathcal{L}_{\text{reg}}$ (9)
$\mathcal{L}_{\text{reg}}$ $=\mathcal{L}_{\text{hard}}:=N^{-1}\sum_{i=1}^{N}\left(\beta\underbracket{\|\mathbf{z}_{i}-\mathtt{detach}(\mathbf{q}_{\iota(\mathbf{z}_{i})})\|^{2}}_{\text{Commitment Loss}}+\underbracket{\|\mathtt{detach}(\mathbf{z}_{i})-\mathbf{q}_{\iota(\mathbf{z}_{i})}\|^{2}}_{\text{Codebook Loss}}\right)$ (10)
where $\mathcal{L}_{\text{main}}$ is the primary task loss (e.g., L2 regression in autoencoding), and $\beta>0$ is a weighting hyperparameter. As defined in in Equation 10, $\mathcal{L}_{\text{hard}}$ consists of two components. The first term (known as the commitment loss) aligns each pre-quantized feature $\mathbf{z}_{i}$ in a batch ($i=1,\dots,N$) with their nearest codebook entry $\mathbf{q}_{\iota(\mathbf{z}_{i})}$. The second term (called the codebook loss) moves each codebook vector $\mathbf{q}_{m}$ toward the centroid of its assigned features whose nearest neighbor is $\mathbf{q}_{m}$ (i.e., $\{\mathbf{z}_{i}:\iota(\mathbf{z}_{i})=m\}$).
It should be noted, however, that the codebook loss in Equation 10 does not inherently prevent code collapse. Some codebook vectors may never serve as the nearest neighbor of any pre-quantized feature and therefore receive no updates. Accordingly, previous studies have resorted to additional workarounds; for example, unused codebook vectors may be reset to the positions of pre-quantized features Dhariwal et al., (2020).
Similarly, smoothed quantization also requires auxiliary regularization to avoid code collapse. A widely adopted option is the normalized perplexity of the mean assignment probability (Dieleman et al.,, 2018; Baevski et al., 2020b, ):
$\mathcal{L}_{\text{reg}}$ $=\mathcal{L}_{\text{ppl}}:=\frac{\exp\left(-\sum_{m}\bar{\pi}_{m}\log\bar{\pi}_{m}\right)}{M}$ (11)
$\bar{\pi}_{m}$ $:=N^{-1}\sum_{i=1}^{N}\pi_{i,m}$ (12)
As noted in §1, this perplexity-based regularization does not promote the onehotness of $\boldsymbol{\pi}_{i}$, failing to distinguish onehot-like samples from uniform or centered ones (Figure 1). Onehotness can instead be induced by annealing the temperature parameter of the Gumbel-softmax sampling ($\tau\to 0$). However, manually scheduling this annealing is empirically challenging. The next section therefore proposes an alternative regularization loss that automatically encourages onehotness within the framework of gradient-based learning, while simultaneously preventing code collapse.
3 Methods
As noted in previous sections, ideal smoothed quantizers $\mathbf{p}\in\Delta^{M-1}$ are distributed near the vertices of the simplex (Figure 1). Moreover, each vertex should have at least some smoothed quantizers in its neighborhood; otherwise, the quantization suffers from code collapse. A simple way to achieve these objectives simultaneously is to impose a loss penalizing the deviation of the KNNs from each vertex (Figure 1):
$\mathcal{L}_{\text{reg}}$ $=\mathcal{L}_{\text{KNN}}:=(MK)^{-1}\sum_{m=1}^{M}\sum_{k=1}^{K}\mathcal{D}(\mathbf{e}_{m},\mathbf{p}^{(m,k)})$ (13)
where $\mathbf{p}^{(m,k)}$ denotes the $k$-th nearest neighbor of the simplex vertex (onehot vector) $\mathbf{e}_{m}$ according to a distance/deviation metric $\mathcal{D}$. Two options for $\mathcal{D}$ are considered in this study: the squared L2 distance, $\|\mathbf{e}_{m}-\mathbf{p}^{(m,k)}\|^{2}$, and cross-entropy, $\log p^{(m,k)}_{m}$.
At first glance, the proposed regularization may appear similar to the commitment and codebook losses used in the gradient-estimation approaches (Equation 10), Both involve nearest neighbors, but the two methods differ in their choice of anchors and neighbors. While commitment and codebook losses take data points as anchors and identify their nearest codebook entries, the proposed method uses codebook entries as anchors and treats data as neighbors. Accordingly, the proposed regularization ensures that every codebook entry receives optimization feedback, whereas commitment/codebook losses may leave some entries untrained if they never become the nearest neighbor of any data point.
A further advantage of the proposed regularization is that Gumbel-softmax sampling is no longer required; smoothed quantizers can be obtained directly as $\mathbf{p}=\boldsymbol{\pi}=\operatorname{softmax}(Q^{\mathsf{T}}\mathbf{z})$. [^2] The exact implementation of $\mathbf{p}$ and $\boldsymbol{\pi}$ involves normalization and rescaling; see Appendix A.1 for details. At the same time, the proposed regularization is fully compatible with Gumbel-softmax sampling; one can simply replace $\mathbf{p}$ (Gumbel-softmax samples) in Equation 13 with $\boldsymbol{\pi}$ (assignment probabilities).
During inference, hard quantization is applied by taking $\operatorname{argmax}_{m}p_{m}$ in the onehot representation. In the following section, both the deterministic and stochastic approaches are evaluated on representative benchmarks.
4 Experiments
The proposed regularization for smoothed vector quantization was benchmarked on two tasks: discrete autoencoding (§4.1) and contrastive learning (§4.2). The Python code used for these experiments will be published when this manuscript is accepted for a journal/conference.
4.1 Discrete Autoencoding
The first benchmark assessed the proposed regularization in the context of discrete autoencoding on the ImageNet dataset (Deng et al.,, 2009). Input images were convolutionally encoded into latent feature maps, whose pixels were then quantized (Esser et al.,, 2021; Fifty et al.,, 2025). The decoder convolutional network reconstructed the input images from these quantized feature maps, and the entire model was trained to minimize the L2 reconstruction loss ($L_{\text{main}}$ in Equation 9). Further details about the network architecture and training setup are provided in Appendix B.1.
Extending prior work (Esser et al.,, 2021; Fifty et al.,, 2025), three different combinations of feature map and codebook sizes were examined. In addition to the previously used settings of $H\times W\times C=16\times 16\times 32$ ($M=1024$) and $64\times 64\times 3$ ($M=8196$), an additional configuration increased the channel dimensionality of the latter to 3232 ($H\times W\times C=64\times 64\times 32,M=8196$).
Each model was trained using four GPUs, and the proposed method computed the $K/4$-nearest neighbors of each simplex vertex per GPU (see §5.3 for further discussion). The number of neighbors was set to $K=2\times 4$ for the spatially smaller feature maps of size $16\times 16\times 32$, and to $K=8\times 4$ for the larger ones. [^3] The value of $K$ was upper-bounded at $8\times 4$ by available computational resources; With a maximum batch size of $64$, the total number of latent pixels was $64\times 64\times 64=8\times 4\times 8196$, allowing only $8\times 4$ neighbors per vertex of $\Delta^{8196}$. This implementation constraint is further discussed in §5.3.
Table 1: Performance of discrete autoencoding on the ImageNet validation set. Reported metrics are codebook usage and reconstruction quality scores: root mean squared error (rMSE), Fréchet Inception Distance (FID), and Inception Score (IS). The proposed method is denoted as “KNN-L2/CE”.
Method Feature Map Size; Codebook Size
$\times16$$\times$32; 1024 $\times64$$\times$3; 8196 $\times64$$\times$32; 8196
Code Use ($\uparrow$) rMSE ($\downarrow$) FID ($\downarrow$) IS ($\uparrow$) Code Use rMSE FID IS Code Use rMSE FID IS
STE
Euclid 4.5% 0.404 124.855 36.608 100.0% 0.167 7.246 402.837 1.7% 0.235 22.019 290.243
Cosine 3.0% 0.381 117.950 41.475 70.9% 0.197 13.972 348.307 2.8% 0.186 12.108 363.862
RE
Euclid 3.1% 0.460 170.300 19.253 78.85% 0.171 10.2086 377.5555 0.5% 0.271 40.575 203.515
Cosine 2.8% 0.423 157.767 24.327 99.5% 0.19 14.672 344.166 4.4% 0.180 10.833 372.889
H-Gmb
PPL 100.0% 0.349 100.286 54.042 100.0% 0.189 19.069 321.164 100.0% 0.163 10.331 384.041
KNN-L2 20.5% 0.354 100.467 56.053 99.9% 0.222 25.002 280.972 23.6% 0.185 10.718 376.354
KNN-CE 100.0% 0.390 110.974 44.817 100.0% 0.226 12.787 349.677 100.0% 0.211 9.343 371.964
S-Gmb
PPL 55.0% 0.826 173.593 9.605 100.0% 0.183 10.278 377.787 48.4% 0.386 45.606 167.142
KNN-L2 100.0% 0.371 81.479 71.946 100.0% 0.201 10.168 373.830 100.0% 0.193 5.727 404.184
KNN-CE 100.0% 0.569 194.736 16.595 100.0% 0.227 13.647 345.019 100.0% 0.218 9.205 366.454
Softmax
PPL 100.0% 1.112 309.603 4.552 81.6% 0.701 37.987 196.723 99.8% 0.725 83.070 77.981
KNN-L2 100.0% 0.366 81.691 71.750 100.0% 0.199 12.644 361.228 100.0% 0.204 7.908 381.654
KNN-CE 100.0% 0.380 106.048 50.750 100.0% 0.219 14.241 343.228 100.0% 0.199 5.267 404.098
Table 3 reports the codebook usage and reconstruction quality scores—including root mean squared error (rMSE), Inception Score (IS; Salimans et al.,, 2016), and Fréchet Inception Distance (FID; Heusel et al.,, 2017)–for the proposed method (“KNN-L2/CE”) and the baselines. The results indicate that reconstruction through softmax-smoothed quantization (i.e., without Gumbel randomness) deteriorated substantially when combined with the perplexity-based regularization (“PPL”; Equation 11). This degradation stems from the mismatch between soft quantization during the training and hard quantization at inference, supporting the argument made in the Introduction that perplexity-based regularization alone does not promote tight smoothing (i.e., vertex-concentrated distributions).
The perplexity-based regularization was only effective when combined with the Gumbel-softmax sampling, and onehot quantization in the forward computation by STE was necessary to ensure full codebook usage (“Hard-Gumbel”). Otherwise, code collapse was not prevented when the channel dimensionality was large ($C=64$; “Soft-Gumbel”).
By contrast, the proposed KNN-based regularization successfully prevented code collapse and achieved near-complete codebook utilization without resorting to Gumbel-softmax sampling. The reconstruction quality was also superior or competitive with the Gumbel-softmax + perplexity approach across all feature map sizes.
The choice of deviation metric (L2 distance vs. cross-entropy, CE) did not result in consistent global superiority, although noticeable code collapse occurred when KNN-L2 regularization was combined with hard Gumbel-softmax sampling, which nonetheless had limited effect on the reconstruction quality.
It is also worth noting that both STE and rotational gradient estimation (RE) suffered from code collapse when the channel dimensionality was large ($C=64$), illustrating the difficulty of achieving robust prevention of code collapse across different settings. The next section further highlights this challenge under a different training objective.
4.2 Contrastive Learning
The second experiment evaluated vector quantization methods within Wav2Vec 2.0 pretraining for speech feature extraction (Baevski et al., 2020b, ). Unlike autoencoding, this pretraining integrates vector quantization directly into the main loss function.
Two codebook configurations were investigated. The first used a single codebook of size 1024. By contrast, the second followed the original work of Baevski et al., 2020b , combining two smaller codebooks—each of size 320—to implement rich code diversity efficiently via product quantization (Jégou et al.,, 2011, see §5.3 for more information). The dimensionality of the codebook vectors was set to 256 for the single-codebook configuration and to 128 for the dual-codebook configuration. The number of nearest neighbors was set to $K=2\times 4$, meaning that two neighbors were identified per codebook entry on each of the four GPUs. In the dual-codebook condition, the two codebooks were treated independently.
Models were trained on the LibriSpeech dataset, combining all training splits (train-clean-100 + train-clean-360 + train-other-500; Panayotov et al.,, 2015). Further details on the network architecture and learning objective are provided in Appendix B.2.
Table 2: Codebook usage in Wav2Vec 2.0 pretraining, evaluated on the LibriSpeech dev-clean split (similar results were observed for the other dev/test splits). The proposed method is denoted as “KNN-L2/CE”.
#Codebooks $\times$ Codebook Size
[…]
Codebook#1 Codebook#2
STE
Euclid 0.2% 0.9% 0.9%
Cosine 0.2% 0.6% 0.6%
RE
Euclid 0.2% 0.6% 0.6%
Cosine 0.2% 0.6% 0.6%
H-Gmb
PPL 0.7% 0.6% 0.6%
KNN-L2 0.2% 0.6% 0.6%
KNN-CE 99.7% 100.0% 100.0%
S-Gmb
PPL 0.3% 0.6% 0.6%
KNN-L2 90.1% 0.6% 0.6%
KNN-CE 100.0% 100.0% 100.0%
Softmax
PPL 0.2% 1.2% 1.2%
KNN-L2 60.4% 100.0% 100.0%
KNN-CE 100.0% 100.0% 100.0%
Table 4 reports codebook usage. The perplexity-based regularization failed to prevent code collapse in both single- and dual-codebook settings. Likewise, the STE and rotational gradient estimation approaches exhibited the same failure.
By contrast, the proposed KNN-based regularization ensured complete code utilization in both conditions when cross-entropy was used as the divergence metric on the simplex. The L2 metric, on the other hand, proved insufficient to prevent code collapse, especially when combined with Gumbel sampling.
5 Discussions
5.1 Summary of Findings & Contributions
This study introduced a simple and unified regularization method that simultaneously tightens smoothed vector quantization and promote effective code utilization. The proposed method successfully prevented code collapse in two representative applications of vector quantization: a middle layer in discrete autoencoding (§4.1) and target construction in contrastive learning (§4.2). This robustness is noteworthy, as prior approaches were effective only in specific settings and remained vulnerable to code collapse in others.
The proposed method is geometrically intuitive and straightforward, yet appears unaddressed in the existing literature. Research on neural vector quantization has traditionally been rooted in variational autoencoding (Kingma and Welling,, 2014), primarily aiming to extend this stochastic framework to discrete variables (Jang et al.,, 2017; van den Oord et al.,, 2017). The issue of code collapse was recognized (or documented) later, and workarounds were developed independently of the quantization methods themselves (Dieleman et al.,, 2018; Baevski et al., 2020b, ; Dhariwal et al.,, 2020).
By contrast, the present work reformulates neural vector quantization as a simple smoothing problem: onehot vectors are approximated by elements of the simplex. Within this perspective, concentrating the approximators near the simplex vertices naturally arises as a desirable property. This reformulation, together with the proposed regularization strategy, represents a key conceptual contribution of the study.
5.2 Alternative Implementations of the Intended Regularization
Alternative regularization strategies could also achieve the intended distribution of smoothed quantizers ($\mathbf{p}$ or $\boldsymbol{\pi}$) around the simplex vertices. For example, one could align smoothed quantizers with a Dirichlet distribution with concentration parameters $\alpha_{1}=\cdots=\alpha_{K}<1.0$ (Figure 2). The probability density of such a distribution is highest at the vertices of the simplex, thus matching the ideal distribution of smoothed quantizers. A possible formalization of this alignment is based on the Kullback-Leibler (KL) divergence between the Dirichlet prior ($\mathbb{P}$) and the distribution inferred from smoothed quantizers ($\mathbb{Q}$).
$\mathcal{D}_{\text{KL}}\left(\mathbb{P}\mid\mathbb{Q}\right):=$ $\int_{\Delta^{M-1}}\mathbb{P}(\mathbf{p})\log\frac{\mathbb{P}(\mathbf{p})}{\mathbb{Q}(\mathbf{p})}d\mathbf{p}$ (14)
However, Equation 14 is difficult to use directly as a regularization loss, due to the complexity of estimating $\mathbb{Q}$ from sample quantizers, $\mathbf{p}$. Although one could constrain $\mathbb{Q}$ as another Dirichlet to make the KL divergence tractable, maximum likelihood estimation of its parameters requires iterative algorithms (e.g., the Newton-Raphson method; Ronning,, 1989; Sklar,, 2014; Wicker et al.,, 2008), complicating and slowing gradient-based optimization in deep learning frameworks. Moreover, this estimation involves digamma and trigamma functions (Sklar,, 2014), whose derivatives can explode when the concentration parameters approach small values—as desired for tight smoothing—during the course of learning.
Figure 2: Dirichlet distributions on the simplex $\Delta^{3-1}$ with concentration parameters $\alpha_{1}=\alpha_{2}=\alpha_{3}=\alpha$, where $\alpha\in\{0.5,1.0,2.0\}$.
A more practical approach is to approximate Equation 14 itself in a tractable manner. For instance, Perez-Cruz, (2008) proposed a KNN-based estimation of the KL divergence that relies solely on samples from the two distributions, using the $k$-th nearest neighbor from $\mathbb{Q}$ for each sample from $\mathbb{P}$. The regularization method proposed in this study can thus be interpreted as minimizing this estimated KL divergence, with the samples from $\mathbb{P}$ constrained to onehot vectors.
5.3 Limitations
A primary limitation of the proposed method is its memory requirement. When training across multiple GPUs, the KNN-based regularization identifies $K$ nearest smoothed quantizers ($\mathbf{p}$ or $\boldsymbol{\pi}$) per simplex vertex on each GPU, rather than finding global neighbors across all GPUs. Consequently, each GPU must have sufficient VRAM to store at least $KM$ latent pixels/frames, where $M$ denotes the codebook size. This requirement can become prohibitive when $M$ is large (i.e., for fine-grained quantization).
One possible workaround is to randomly select a subset of simplex vertices when computing the regularization loss, rather than using all vertices in a single iteration. In expectation, this achieves the same effect as the original implementation, although its empirical effectiveness remains to be assessed in future studies.
Additionally, fine-grained quantization can be achieved more efficiently using smaller $G$ codebooks in combination (product quantization; Jégou et al.,, 2011), as explored in the Wav2Vec 2.0 pretraining (§4.2). This approach represents $\prod_{g=1}^{G}M_{g}$ distinct quantized vectors while requiring only $K\sum_{g=1}^{G}M_{g}$ smoothed quantizers per GPU. Leveraging these strategies, the proposed method can overcome its limitation and become applicable to real-world scenarios.
Acknowledgments
This study was supported by JSPS Grant-in-Aid for Scientific Research A (JP24H00774), B (JP22H03914), and C (JP24K15087); JST AIP Accelerated Program (JPMJCR25U6) and Core Research for Evolutional Science and Technology (JPMJCR22P5); and Kayamori Foundation of Informational Science Advancement (K35XXVIII620). The author also gratefully acknowledges the support of the Academic Center for Computing and Media Studies, Kyoto University, regarding the use of their supercomputer system.
Appendix A Implementation of the Quantization Methods
This section provides details on the implementation of the quantization methods.
A.1 Smoothed Quantization
Smoothed quantizers $\mathbf{p}$ were computed as $\mathbf{p}=\operatorname{softmax}(\hat{Q}^{\mathsf{T}}\hat{\mathbf{z}}/\tau)$. In other words, the codebook vectors $(\mathbf{q}_{1},\dots,\mathbf{q}_{M})=Q$ and the feature vectors $\mathbf{z}$ were first L2-normalized, and their product (i.e., cosine similarity) was rescaled by a learnable temperature $\tau$. This temperature was shared across the codebook so that all the logits had the same amplitude. Assignment probabilities $\mathcal{\pi}$ for Gumbel-softmax sampling were computed in the same way. [^4] Previous studies manually annealed the temperature $\tau$ instead of treating it as a learnable parameter (Baevski et al., 2020b, ). This approach was also tested in the experiments here but did not yield improvements over the learnable temperature.
A.2 Hard Quantization
The weight $\beta$ on the commitment loss in Equation 10 was set to 1.0, based on the previous observations that its value does not significantly affect learning outcomes within the range 0.1–2.0. (van den Oord et al.,, 2017).
Appendix B Details of the Experiments
B.1 Discrete Autoencoding
This section provides implementation details for the autoencoding experiment described in §4.1.
The network architecture followed prior work on discrete autoencoding of ImageNet (Esser et al.,, 2021; Fifty et al.,, 2025). Input images were center-cropped to $H\times W\times C=256\times 256\times 3$. The encoder first expanded the channel dimensionality of the input images from 3 to 256 by convolution, and then progressively downsampled them through a series of strided convolutions (see Table 3 for the spatial and channel sizes at each layer). Each downsampling layer was followed by a residual block (He et al.,, 2016). The decoder reconstructed the input images by upsampling the latent feature maps with a sequence of interpolations and residual blocks.
To improve memory efficiency—particularly important for the proposed KNN-based regularization—all but the input and output layers were implemented as depthwise separable convolutions (Chollet,, 2017). All convolutional kernels had size $3\times 3$.
Training employed the AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.99)$ and a weight decay coefficient of $10^{-4}$, except for Euclidean-based STE/rotational hard quantization, where weight decay was set to zero. The learning rate was linearly warmed up from $0.0$ to $\rho_{\mathrm{max}}$, and subsequently annealed to $0.5\rho_{\mathrm{max}}$ by cosine scheduling. The maximum learning rate $\rho_{\mathrm{max}}$ was set to $5\times 10^{-5}$ for Euclidean-based STE/rotational hard quantization, and to $10^{-4}$ for all other configurations (Fifty et al.,, 2025).
Table 3: Hyperparameters for discrete autoencoding.
Feature Map Size $16\times 16\times 32$ $\times 64\times 3 64×64$$\times 64\times 32$
Codebook Size 1024 8196 8196
Latent Channels $128\to 128\to 64\to 64\to 32\to 32$ $128\to 64\to 32\to 3$ $128\to 64\to 32\to 32$
Height & Width $256\to 128\to 64\to 32\to 16\to 16$ $256\to 128\to 64\to 64$ $256\to 128\to 64\to 64$
Batch Size 64 64 64
Training Epochs 25 20 20
Warmup Iterations 16,000 16,000 16,000
Inception Score (IS; Salimans et al.,, 2016) and Fréchet Inception Distance (FID; Heusel et al.,, 2017) were estimated using the ImageNet-pretrained Inception V3 provided in torchvision.
B.2 Wav2Vec 2.0
This section provides details for the Wav2Vec 2.0 pretraining discussed in §4.2.
[…]
$\mathcal{L}_{\text{main}}=$ $-\log\frac{\exp(\hat{\mathbf{y}}_{t}\hat{\mathbf{q}}_{t}^{\mathsf{T}}/\mathcal{T})}{\sum_{\hat{\mathbf{q}}\sim\mathcal{Q}}\exp(\hat{\mathbf{y}}_{t}\hat{\mathbf{q}}^{\mathsf{T}}/\mathcal{T})}$ (15)
where $\hat{\cdot}$ denotes the L2-normalization of vectors (i.e., measuring the cosine similarity), and $\mathcal{T}:=0.1$ is the temperature parameter. For each masked vector, a set of distractors $\mathbf{\tilde{q}}$ was sampled from the other quantized vectors according to a distribution $\mathcal{Q}$.
Both the convolutional encoder and Transformer were implemented using the publicly available code in torchaudio, and only the quantization module was implemented from scratch.
Input waveforms were randomly cropped to the length of 250k samples. Both stages employed the AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.99)$ and zero weight decay. The learning rate was warmed up from $0.0$ to $5.0\times 10^{-4}$, and then annealed to $5.0\times 10^{-6}$ by cosine scheduling.
Table 4: Hyperparameters for Wav2Vec 2.0.
Input Frequency 16kHz
Latent Frequency 50Hz
Codebook Dimensionality 256
Codebook Size 1024
CNN
Latent Channels 512
[…]
Transformer
# Layers 12
Model Dimensionality 768
# Heads 8
Feed-Forward Dimensionality 4096
Dropout Rate 0.1
Layer Drop 0.05
[…]
Batch Size 128 64
Training Epochs 128 20
Warmup Iterations 32,000 10,000
When the single-codebook condition was first examined, training was run for 128 epochs, following the original schedule (Baevski et al., 2020b, ). However, since convergence occurred rapidly, the number of epochs was reduced in the dual-codebook condition to improve time efficiency.

### v2

@@
 a codebook).
-The resulting discrete representations are easier to manipulate and interpret than the original continuous forms, and have proven effective across diverse applications, including image generation (Esser et al.,, 2021; Ramesh et al.,, 2021; Rombach et al.,, 2022; Yu et al.,, 2022), speech recognition (Baevski et al., 2020a, ; Baevski et al., 2020b, ), and music generation (Hadjeres and Crestel,, 2020; Dhariwal et al.,, 2020).
+The resulting discrete representations are easier to manipulate and interpret than the original continuous forms, and have proven effective across diverse applications, including image generation (Esser et al.,, 2021; Ramesh et al.,, 2021; Rombach et al.,, 2022; Yu et al., 2022b, ), speech recognition (Baevski et al., 2020a, ; Baevski et al., 2020b, ), and music generation (Hadjeres and Crestel,, 2020; Dhariwal et al.,, 2020).
 When integrated into deep neural networks, however, vector quantization introduces a fundamental challenge: quantization is a non-differentiable operation that blocks gradient backpropagation (van den Oord et al.,, 2017).
@@
 It should be noted, however, that the codebook loss in Equation 10 does not inherently prevent code collapse.
-Some codebook vectors may never serve as the nearest neighbor of any pre-quantized feature and therefore receive no updates.
+Some codebook vectors may never serve as the nearest neighbor of any pre-quantized feature and therefore receive no updates (Zhu et al.,, 2025).
 Accordingly, previous studies have resorted to additional workarounds; for example, unused codebook vectors may be reset to the positions of pre-quantized features Dhariwal et al., (2020).
+More recently, Zhu et al., (2025) proposed another remedy called SimVQ, which reparameterizes the codebook $Q$ as the product of a randomly frozen matrix $Q^{\prime}\in\mathbb{R}^{M\times D}$ and a learnable matrix $W\in\mathbb{R}^{D\times D}$ (i.e., $Q=Q^{\prime}W$).
+In this formulation, all codebook vectors share learnable parameters with one another ($\mathbf{q}_{m}=q^{\prime}_{m,1}\mathbf{w}_{1}+\cdots+q^{\prime}_{m,D}\mathbf{w}_{D}$), so updates to one vector propagate to the others.
 Similarly, smoothed quantization also requires auxiliary regularization to avoid code collapse.
@@
 The proposed regularization for smoothed vector quantization was benchmarked on two tasks: discrete autoencoding (§4.1) and contrastive learning (§4.2).
-The Python code used for these experiments will be published when this manuscript is accepted for a journal/conference.
+The Python code used for these experiments is available as supplementary material.
 4.1 Discrete Autoencoding
@@
 Extending prior work (Esser et al.,, 2021; Fifty et al.,, 2025), three different combinations of feature map and codebook sizes were examined.
-In addition to the previously used settings of $H\times W\times C=16\times 16\times 32$ ($M=1024$) and $64\times 64\times 3$ ($M=8196$), an additional configuration increased the channel dimensionality of the latter to 3232 ($H\times W\times C=64\times 64\times 32,M=8196$).
+In addition to the previously used settings of $H\times W\times C=16\times 16\times 32$ ($M=1024$) and $64\times 64\times 3$ ($M=8196$), an additional configuration increased the channel dimensionality of the latter to $32$ ($H\times W\times C=64\times 64\times 32,M=8196$).
 Each model was trained using four GPUs, and the proposed method computed the $K/4$-nearest neighbors of each simplex vertex per GPU (see §5.3 for further discussion).
-The number of neighbors was set to $K=2\times 4$ for the spatially smaller feature maps of size $16\times 16\times 32$, and to $K=8\times 4$ for the larger ones.
+The number of neighbors was set to $K/4=8$ for the L2 distance and $K/4=1$ for the cross-entropy otherwise specified.
 [^3] The value of $K$ was upper-bounded at $8\times 4$ by available computational resources; With a maximum batch size of $64$, the total number of latent pixels was $64\times 64\times 64=8\times 4\times 8196$, allowing only $8\times 4$ neighbors per vertex of $\Delta^{8196}$.
 This implementation constraint is further discussed in §5.3.
+The weight $\beta$ on the commitment loss (Equation 10) in STE (including SimVQ) and rotational gradient estimation was set to 1.0, following Fifty et al., (2025) and Zhu et al., (2025).
 Table 1: Performance of discrete autoencoding on the ImageNet validation set.
@@
 The proposed method is denoted as “KNN-L2/CE”.
-Method Feature Map Size; Codebook Size
+Best scores across all methods are highlighted in boldface, while underlined values indicate the best-performing number of nearest neighbors among $K/4\in\{1,2,4,8\}$.
+Feature Map Size; Codebook Size
 $\times16$$\times$32; 1024 $\times64$$\times$3; 8196 $\times64$$\times$32; 8196
-Code Use ($\uparrow$) rMSE ($\downarrow$) FID ($\downarrow$) IS ($\uparrow$) Code Use rMSE FID IS Code Use rMSE FID IS
+Method $K/4$ Code Use ($\uparrow$) rMSE ($\downarrow$) FID ($\downarrow$) IS ($\uparrow$) Code Use rMSE FID IS Code Use rMSE FID IS
 STE
-Euclid 4.5% 0.404 124.855 36.608 100.0% 0.167 7.246 402.837 1.7% 0.235 22.019 290.243
-Cosine 3.0% 0.381 117.950 41.475 70.9% 0.197 13.972 348.307 2.8% 0.186 12.108 363.862
+Euclid 4.5% 0.404 124.86 36.61 100.0% 0.167 7.25 402.84 1.7% 0.235 22.02 290.24
+Cosine 3.0% 0.381 117.95 41.48 70.9% 0.197 13.97 348.31 2.8% 0.186 12.11 363.86
+SimVQ 100.0% 0.340 87.26 71.33 100.0% 0.170 7.44 400.41 100.0% 0.148 3.97 436.29
 RE
-Euclid 3.1% 0.460 170.300 19.253 78.85% 0.171 10.2086 377.5555 0.5% 0.271 40.575 203.515
-Cosine 2.8% 0.423 157.767 24.327 99.5% 0.19 14.672 344.166 4.4% 0.180 10.833 372.889
+Euclid 3.1% 0.460 170.30 19.25 78.85% 0.171 10.21 377.56 0.5% 0.271 40.58 203.52
+Cosine 2.8% 0.423 157.77 24.33 99.5% 0.194 14.67 344.17 4.4% 0.180 10.83 372.89
 H-Gmb
-PPL 100.0% 0.349 100.286 54.042 100.0% 0.189 19.069 321.164 100.0% 0.163 10.331 384.041
-KNN-L2 20.5% 0.354 100.467 56.053 99.9% 0.222 25.002 280.972 23.6% 0.185 10.718 376.354
-KNN-CE 100.0% 0.390 110.974 44.817 100.0% 0.226 12.787 349.677 100.0% 0.211 9.343 371.964
+PPL 100.0% 0.349 100.29 54.04 100.0% 0.189 19.07 321.16 100.0% 0.163 10.33 384.04
+KNN-L2 8 52.1% 0.344 98.98 59.66 99.9% 0.222 25.00 280.97 23.6% 0.185 10.72 376.35
+KNN-CE 1 100.0% 0.368 86.49 68.27 100.0% 0.226 14.27 342.64 100.0% 0.173 3.17 435.04
 S-Gmb
-PPL 55.0% 0.826 173.593 9.605 100.0% 0.183 10.278 377.787 48.4% 0.386 45.606 167.142
-KNN-L2 100.0% 0.371 81.479 71.946 100.0% 0.201 10.168 373.830 100.0% 0.193 5.727 404.184
-KNN-CE 100.0% 0.569 194.736 16.595 100.0% 0.227 13.647 345.019 100.0% 0.218 9.205 366.454
+PPL 55.0% 0.826 173.59 9.61 100.0% 0.183 10.28 377.79 48.4% 0.386 45.61 167.14
+KNN-L2 8 96.3% 0.358 74.00 85.36 100.0% 0.201 10.17 373.83 100.0% 0.193 5.73 404.18
+KNN-CE 1 100.0% 0.569 194.74 16.60 100.0% 0.233 15.41 329.78 100.0% 0.186 3.83 425.33
 Softmax
-PPL 100.0% 1.112 309.603 4.552 81.6% 0.701 37.987 196.723 99.8% 0.725 83.070 77.981
-KNN-L2 100.0% 0.366 81.691 71.750 100.0% 0.199 12.644 361.228 100.0% 0.204 7.908 381.654
-KNN-CE 100.0% 0.380 106.048 50.750 100.0% 0.219 14.241 343.228 100.0% 0.199 5.267 404.098
+PPL 100.0% 1.112 309.60 4.55 81.6% 0.701 37.99 196.72 99.8% 0.725 83.07 77.98
+KNN-L2 1 100.0% 0.358 76.29 81.61 100.0% 0.224 12.08 354.36 100.0% 0.387 32.20 210.61
+2 100.0% 0.366 81.69 71.75 100.0% 0.226 12.04 348.48 100.0% 0.175 3.62 427.42
+4 100.0% 0.379 87.05 62.37 100.0% 0.205 9.01 383.56 100.0% 0.196 5.64 403.06
+8 99.9% 0.343 73.72 88.53 100.0% 0.199 12.64 361.23 100.0% 0.204 7.91 381.65
+KNN-CE 1 100.0% 0.366 79.80 74.69 100.0% 0.225 14.08 345.56 100.0% 0.175 2.81 437.72
+2 100.0% 0.380 106.05 50.75 100.0% 0.227 14.34 338.37 100.0% 0.187 4.16 418.57
+4 100.0% 0.748 255.59 9.03 100.0% 0.228 16.73 326.78 100.0% 0.204 6.98 387.88
+8 100.0% 1.240 467.86 3.01 100.0% 0.219 14.24 343.23 100.0% 0.199 5.27 404.10
 Table 1 reports the codebook usage and reconstruction quality scores—including root mean squared error (rMSE), Inception Score (IS; Salimans et al.,, 2016), and Fréchet Inception Distance (FID; Heusel et al.,, 2017)–for the proposed method (“KNN-L2/CE”) and the baselines.
-The results indicate that reconstruction through softmax-smoothed quantization (i.e., without Gumbel randomness) deteriorated substantially when combined with the perplexity-based regularization (“PPL”; Equation 11).
-This degradation stems from the mismatch between soft quantization during the training and hard quantization at inference, supporting the argument made in the Introduction that perplexity-based regularization alone does not promote tight smoothing (i.e., vertex-concentrated distributions).
+Reconstruction through softmax-smoothed quantization (i.e., without Gumbel randomness) deteriorated substantially when combined with the perplexity-based regularization (“PPL”; Equation 11).
+This degradation stems from the mismatch between soft quantization during the training and hard quantization at inference; the smoothed quantizers deviated from the simplex vertices, as reflected in their high individual perplexity, $\exp(-\sum_{m=1}^{M}p_{m}\log p_{m})$ (Table 2).
+These observations support the argument made in the Introduction that perplexity-based regularization alone does not promote tight smoothing.
+Table 2: Tightness of softmax-based smoothing (without Gumbel sampling), measured by the individual perplexity of smoothed quantizers, $\exp(-\sum_{m=1}^{M}p_{m}\log p_{m})$.
+Reported values are the 75th, 90th, and 99th percentiles, as well as the maximum, computed across all feature-map pixels in the ImageNet validation set.
+Feature Map Size; Codebook Size
+$\times16$$\times$32; 1024 $\times64$$\times$3; 8192 $\times64$$\times$32; 8192
+Method $K/4$ 75% 90% 99% Max 75% 90% 99% Max 75% 90% 99% Max
+PPL 910.40 933.12 963.29 995.91 7559.89 7561.99 7563.10 7565.02 7230.60 7267.06 7315.62 7424.16
+KNN-L2 1 1.17 1.59 2.36 6.26 1.00 1.00 1.00 4.00 2.24 40.45 86.53 249.64
+2 1.00 1.08 1.86 27.37 1.00 1.00 1.06 4.65 1.14 1.55 2.33 7.92
+4 1.00 1.06 1.86 62.99 1.00 1.00 1.00 4.61 1.00 1.06 1.80 20.42
+8 1.00 1.10 1.90 120.68 3.33 3.87 5.22 12.03 1.00 1.05 1.79 107.70
+KNN-CE 1 1.18 1.60 2.38 6.28 3.96 4.51 5.81 11.96 1.09 1.45 2.17 5.94
+2 1.23 1.72 2.73 8.07 4.04 4.66 6.16 11.77 1.06 1.38 2.08 6.15
+4 1.42 1.78 2.38 5.45 4.16 4.86 6.57 13.04 1.08 1.44 2.16 6.58
+8 2.38 2.80 3.71 7.78 4.46 5.31 7.39 13.81 1.29 1.75 2.65 9.26
 The perplexity-based regularization was only effective when combined with the Gumbel-softmax sampling, and onehot quantization in the forward computation by STE was necessary to ensure full codebook usage (“Hard-Gumbel”).
-Otherwise, code collapse was not prevented when the channel dimensionality was large ($C=64$; “Soft-Gumbel”).
+Otherwise, code collapse was not prevented when the channel dimensionality was large ($C=32$; “Soft-Gumbel”).
 By contrast, the proposed KNN-based regularization successfully prevented code collapse and achieved near-complete codebook utilization without resorting to Gumbel-softmax sampling.
@@
 cross-entropy, CE) did not result in consistent global superiority, although noticeable code collapse occurred when KNN-L2 regularization was combined with hard Gumbel-softmax sampling, which nonetheless had limited effect on the reconstruction quality.
-It is also worth noting that both STE and rotational gradient estimation (RE) suffered from code collapse when the channel dimensionality was large ($C=64$), illustrating the difficulty of achieving robust prevention of code collapse across different settings.
-The next section further highlights this challenge under a different training objective.
+In terms of computational efficiency, however, cross-entropy proved more favorable.
+Notably, its performance was highest when the number of neighbors was minimal ($K/4=1$) across all feature map sizes (indicated by the underlined scores in Table 1); increasing $K$ did not yield further improvements.
+Although most smoothed quantizers remained unregularized under this setting, they still achieved low individual perplexities after the training (Table 2), indicating a tight approximation of quantization.
+By contrast, minimizing only the L2 distance between the simplex vertices and their single nearest neighbor per GPU led to performance degradation for the feature map of size $64\times 64\times 8192$.
+Since requiring more neighbors necessitates a larger batch size (see §5.3 for details), cross-entropy emerges as a more scalable option.
+Finally, both STE and rotational gradient estimation (RE) exhibited severe code collapse when the channel dimensionality was large ($C=32$), reaffirming prior findings that reducing channel dimensionality is necessary for stable training (Yu et al., 2022a, ; Yu et al.,, 2024; Mentzer et al.,, 2024).
+By contrast, SimVQ maintained full codebook usage across all settings and achieved strong reconstruction quality even at high dimensionality.
+However, the next section presents a case study in which all gradient-estimation approaches—including SimVQ—encounter code collapse, highlighting the difficulty of achieving robust prevention of code collapse across different settings.
 4.2 Contrastive Learning
@@
 The dimensionality of the codebook vectors was set to 256 for the single-codebook configuration and to 128 for the dual-codebook configuration.
-The number of nearest neighbors was set to $K=2\times 4$, meaning that two neighbors were identified per codebook entry on each of the four GPUs.
-In the dual-codebook condition, the two codebooks were treated independently.
-Models were trained on the LibriSpeech dataset, combining all training splits (train-clean-100 + train-clean-360 + train-other-500; Panayotov et al.,, 2015).
-Further details on the network architecture and learning objective are provided in Appendix B.2.
+The weight $\beta$ on the commitment loss in both STE (including SimVQ) and rotational gradient estimation was set to 1.0 (Fifty et al.,, 2025; Zhu et al.,, 2025).
 Table 3: Codebook usage in Wav2Vec 2.0 pretraining, evaluated on the LibriSpeech dev-clean split (similar results were observed for the other dev/test splits).
@@
 […]
-Codebook#1 Codebook#2
+Method $K/4$ Codebook#1 Codebook#2
 STE
-Euclid 0.2% 0.9% 0.9%
+Euclid 0.8% 0.9% 0.9%
 Cosine 0.2% 0.6% 0.6%
+SimVQ 0.2% 0.6% 0.6%
 RE
-Euclid 0.2% 0.6% 0.6%
+Euclid 2.5% 0.6% 0.6%
 Cosine 0.2% 0.6% 0.6%
@@
 PPL 0.7% 0.6% 0.6%
-KNN-L2 0.2% 0.6% 0.6%
-KNN-CE 99.7% 100.0% 100.0%
+KNN-L2 2 0.2% 0.6% 0.6%
+KNN-CE 2 99.7% 100.0% 100.0%
 S-Gmb
 PPL 0.3% 0.6% 0.6%
-KNN-L2 90.1% 0.6% 0.6%
-KNN-CE 100.0% 100.0% 100.0%
+KNN-L2 2 90.1% 0.6% 0.6%
+KNN-CE 2 100.0% 100.0% 100.0%
 Softmax
 PPL 0.2% 1.2% 1.2%
-KNN-L2 60.4% 100.0% 100.0%
-KNN-CE 100.0% 100.0% 100.0%
+KNN-L2 1 82.4% 100.0% 100.0%
+2 60.4% 100.0% 100.0%
+KNN-CE 1 99.5% 100.0% 100.0%
+2 100.0% 100.0% 100.0%
+Models were trained on the LibriSpeech dataset, combining all training splits (train-clean-100 + train-clean-360 + train-other-500; Panayotov et al.,, 2015).
+Further details on the network architecture and learning objective are provided in Appendix B.2.
 Table 3 reports codebook usage.
@@
 Likewise, the STE and rotational gradient estimation approaches exhibited the same failure.
-By contrast, the proposed KNN-based regularization ensured complete code utilization in both conditions when cross-entropy was used as the divergence metric on the simplex.
+Remarkably, SimVQ—despite achieving full codebook usage in the discrete autoencoding experiments—offered no observable benefit in this learning paradigm.
+By contrast, the proposed KNN-based regularization ensured (near-)complete code utilization in both conditions when cross-entropy was used as the divergence metric on the simplex.
+As in the discrete autoencoding experiments, a single neighbor per GPU was sufficient to obtain this effect.
 The L2 metric, on the other hand, proved insufficient to prevent code collapse, especially when combined with Gumbel sampling.
@@
 Consequently, each GPU must have sufficient VRAM to store at least $KM$ latent pixels/frames, where $M$ denotes the codebook size.
-This requirement can become prohibitive when $M$ is large (i.e., for fine-grained quantization).
+This requirement can become prohibitive when $M$ is large (i.e., for fine-grained quantization), although empirically, a single neighbor per GPU appeared sufficient to prevent code collapse when cross-entropy was used as the divergence metric.
 One possible workaround is to randomly select a subset of simplex vertices when computing the regularization loss, rather than using all vertices in a single iteration.
@@
 Acknowledgments
-This study was supported by JSPS Grant-in-Aid for Scientific Research A (JP24H00774), B (JP22H03914), and C (JP24K15087); JST AIP Accelerated Program (JPMJCR25U6) and Core Research for Evolutional Science and Technology (JPMJCR22P5); and Kayamori Foundation of Informational Science Advancement (K35XXVIII620).
-The author also gratefully acknowledges the support of the Academic Center for Computing and Media Studies, Kyoto University, regarding the use of their supercomputer system.
+Hidden for double-blind review.
 Appendix A Implementation of the Quantization Methods
@@
 A.1 Smoothed Quantization
-Smoothed quantizers $\mathbf{p}$ were computed as $\mathbf{p}=\operatorname{softmax}(\hat{Q}^{\mathsf{T}}\hat{\mathbf{z}}/\tau)$.
-In other words, the codebook vectors $(\mathbf{q}_{1},\dots,\mathbf{q}_{M})=Q$ and the feature vectors $\mathbf{z}$ were first L2-normalized, and their product (i.e., cosine similarity) was rescaled by a learnable temperature $\tau$.
-This temperature was shared across the codebook so that all the logits had the same amplitude.
-Assignment probabilities $\mathcal{\pi}$ for Gumbel-softmax sampling were computed in the same way.
-[^4] Previous studies manually annealed the temperature $\tau$ instead of treating it as a learnable parameter (Baevski et al., 2020b, ).
-This approach was also tested in the experiments here but did not yield improvements over the learnable temperature.
+[…]
 A.2 Hard Quantization
-The weight $\beta$ on the commitment loss in Equation 10 was set to 1.0, based on the previous observations that its value does not significantly affect learning outcomes within the range 0.1–2.0.
+The weight $\beta$ on the commitment loss in Equation 10 was set to 1.0 (Fifty et al.,, 2025; Zhu et al.,, 2025), based on the previous observations that its value does not significantly affect learning outcomes within the range 0.1–2.0.
 (van den Oord et al.,, 2017).
@@
 Both the convolutional encoder and Transformer were implemented using the publicly available code in torchaudio, and only the quantization module was implemented from scratch.
+All components are provided in the supplementary material.
 Input waveforms were randomly cropped to the length of 250k samples.

### v3

@@
 The resulting discrete representations are easier to manipulate and interpret than the original continuous forms, and have proven effective across diverse applications, including image generation (Esser et al.,, 2021; Ramesh et al.,, 2021; Rombach et al.,, 2022; Yu et al., 2022b, ), speech recognition (Baevski et al., 2020a, ; Baevski et al., 2020b, ), and music generation (Hadjeres and Crestel,, 2020; Dhariwal et al.,, 2020).
-When integrated into deep neural networks, however, vector quantization introduces a fundamental challenge: quantization is a non-differentiable operation that blocks gradient backpropagation (van den Oord et al.,, 2017).
-Accordingly, some approximation is required to enable learning through quantization.
-One effective workaround is to smooth vector quantization (Jang et al.,, 2017).
-The selection of a codebook vector can be expressed as multiplying the codebook matrix by its corresponding onehot vector $(0,\dots,1,\dots,0)^{\mathsf{T}}$, whose nonzero entry indexes the chosen vector.
-Smoothed vector quantization relaxes this onehot vector to lie within the simplex $\Delta^{M-1}:=\{(p_{1},\dots,p_{M})\mid\sum_{m=1}^{M}p_{m}=1\}$, where $M$ denotes the number of codebook vectors.
-Consequently, differentiable mappings (e.g., softmax) become available and can be incorporated into neural networks.
 […]
@@
 (1) The proposed $K$-nearest neighbor (KNN) distance minimization ($K=8$) favors the desired vertex-concentrated distribution while also preventing code collapse.
+When integrated into deep neural networks, however, vector quantization introduces a fundamental challenge: quantization is a non-differentiable operation that blocks gradient backpropagation (van den Oord et al.,, 2017).
+Accordingly, some approximation is required to enable learning through quantization.
+One effective workaround is to smooth vector quantization (Jang et al.,, 2017).
+The selection of a codebook vector can be expressed as multiplying the codebook matrix by its corresponding onehot vector $(0,\dots,1,\dots,0)^{\mathsf{T}}$, whose nonzero entry indexes the chosen vector.
+Smoothed vector quantization relaxes this onehot vector to lie within the simplex $\Delta^{M-1}:=\{(p_{1},\dots,p_{M})\mid\sum_{m=1}^{M}p_{m}=1\}$, where $M$ denotes the number of codebook vectors.
+Consequently, differentiable mappings (e.g., softmax) become available and can be incorporated into neural networks.
 To successfully approximate onehot quantizers, smoothed quantizers must be distributed around the vertices of the simplex (orange “o” in Figure 1).
@@
 $\mathcal{L}_{\text{total}}$ $=\mathcal{L}_{\text{main}}+\mathcal{L}_{\text{reg}}$ (9)
-$\mathcal{L}_{\text{reg}}$ $=\mathcal{L}_{\text{hard}}:=N^{-1}\sum_{i=1}^{N}\left(\beta\underbracket{\|\mathbf{z}_{i}-\mathtt{detach}(\mathbf{q}_{\iota(\mathbf{z}_{i})})\|^{2}}_{\text{Commitment Loss}}+\underbracket{\|\mathtt{detach}(\mathbf{z}_{i})-\mathbf{q}_{\iota(\mathbf{z}_{i})}\|^{2}}_{\text{Codebook Loss}}\right)$ (10)
+$\mathcal{L}_{\text{reg}}$ $=\mathcal{L}_{\text{hard}}$
+$:=N^{-1}\sum_{i=1}^{N}\left(\beta\underbracket{\|\mathbf{z}_{i}-\mathtt{detach}(\mathbf{q}_{\iota(\mathbf{z}_{i})})\|^{2}}_{\text{Commitment Loss}}\right$.
+$\qquad\qquad\qquad\left.+\underbracket{\|\mathtt{detach}(\mathbf{z}_{i})-\mathbf{q}_{\iota(\mathbf{z}_{i})}\|^{2}}_{\text{Codebook Loss}}\right)$ (10)
 where $\mathcal{L}_{\text{main}}$ is the primary task loss (e.g., L2 regression in autoencoding), and $\beta>0$ is a weighting hyperparameter.
@@
 Some codebook vectors may never serve as the nearest neighbor of any pre-quantized feature and therefore receive no updates (Zhu et al.,, 2025).
-Accordingly, previous studies have resorted to additional workarounds; for example, unused codebook vectors may be reset to the positions of pre-quantized features Dhariwal et al., (2020).
+Accordingly, previous studies have resorted to additional workarounds; for example, unused codebook vectors may be reset to the positions of pre-quantized features (Dhariwal et al.,, 2020).
 More recently, Zhu et al., (2025) proposed another remedy called SimVQ, which reparameterizes the codebook $Q$ as the product of a randomly frozen matrix $Q^{\prime}\in\mathbb{R}^{M\times D}$ and a learnable matrix $W\in\mathbb{R}^{D\times D}$ (i.e., $Q=Q^{\prime}W$).
@@
 The proposed regularization for smoothed vector quantization was benchmarked on two tasks: discrete autoencoding (§4.1) and contrastive learning (§4.2).
-The Python code used for these experiments is available as supplementary material.
+The Python code used for these experiments will be published once this study is accepted.
 4.1 Discrete Autoencoding
 The first benchmark assessed the proposed regularization in the context of discrete autoencoding on the ImageNet dataset (Deng et al.,, 2009).
-Input images were convolutionally encoded into latent feature maps, whose pixels were then quantized (Esser et al.,, 2021; Fifty et al.,, 2025).
-The decoder convolutional network reconstructed the input images from these quantized feature maps, and the entire model was trained to minimize the L2 reconstruction loss ($L_{\text{main}}$ in Equation 9).
-Further details about the network architecture and training setup are provided in Appendix B.1.
-Extending prior work (Esser et al.,, 2021; Fifty et al.,, 2025), three different combinations of feature map and codebook sizes were examined.
-In addition to the previously used settings of $H\times W\times C=16\times 16\times 32$ ($M=1024$) and $64\times 64\times 3$ ($M=8196$), an additional configuration increased the channel dimensionality of the latter to $32$ ($H\times W\times C=64\times 64\times 32,M=8196$).
-Each model was trained using four GPUs, and the proposed method computed the $K/4$-nearest neighbors of each simplex vertex per GPU (see §5.3 for further discussion).
-The number of neighbors was set to $K/4=8$ for the L2 distance and $K/4=1$ for the cross-entropy otherwise specified.
+Input images were convolutionally encoded into latent feature maps, whose pixels were subsequently quantized (Esser et al.,, 2021; Fifty et al.,, 2025).
+A decoder convolutional network reconstructed the input images from these quantized feature maps, and the entire model was trained to minimize the L2 reconstruction loss ($L_{\text{main}}$ in Equation 9).
+Further details of the network architecture and training setup are provided in Appendix B.1.
+Extending prior work (Esser et al.,, 2021; Fifty et al.,, 2025), four combinations of feature map and codebook sizes were examined.
+In addition to the previously studied configurations of $H\times W\times C=16\times 16\times 32$ ($M=1024$) and $64\times 64\times 3$ ($M=8196$), the channel dimensionality of the latter was further increased to $C=32$ and $C=2048$, yielding feature maps of size $64\times 64\times 32$ and $64\times 64\times 2048$ (both with $M=8196$).
+Each model was trained using four GPUs in parallel.
+Accordingly, the proposed method computed the $K/4$-nearest neighbors of each simplex vertex independently on each GPU (see §5.3 for further discussion).
+The number of neighbors was chosen from $K/4\in\{1,2,4,8\}$.
 [^3] The value of $K$ was upper-bounded at $8\times 4$ by available computational resources; With a maximum batch size of $64$, the total number of latent pixels was $64\times 64\times 64=8\times 4\times 8196$, allowing only $8\times 4$ neighbors per vertex of $\Delta^{8196}$.
 This implementation constraint is further discussed in §5.3.
-The weight $\beta$ on the commitment loss (Equation 10) in STE (including SimVQ) and rotational gradient estimation was set to 1.0, following Fifty et al., (2025) and Zhu et al., (2025).
+The weight $\beta$ on the commitment loss (Equation 10) for STE-based methods (including SimVQ) and rotational gradient estimation was set to 1.0, following Fifty et al., (2025) and Zhu et al., (2025).
 Table 1: Performance of discrete autoencoding on the ImageNet validation set.
@@
 Feature Map Size; Codebook Size
-$\times16$$\times$32; 1024 $\times64$$\times$3; 8196 $\times64$$\times$32; 8196
-Method $K/4$ Code Use ($\uparrow$) rMSE ($\downarrow$) FID ($\downarrow$) IS ($\uparrow$) Code Use rMSE FID IS Code Use rMSE FID IS
-STE
-Euclid 4.5% 0.404 124.86 36.61 100.0% 0.167 7.25 402.84 1.7% 0.235 22.02 290.24
-Cosine 3.0% 0.381 117.95 41.48 70.9% 0.197 13.97 348.31 2.8% 0.186 12.11 363.86
-SimVQ 100.0% 0.340 87.26 71.33 100.0% 0.170 7.44 400.41 100.0% 0.148 3.97 436.29
-RE
-Euclid 3.1% 0.460 170.30 19.25 78.85% 0.171 10.21 377.56 0.5% 0.271 40.58 203.52
-Cosine 2.8% 0.423 157.77 24.33 99.5% 0.194 14.67 344.17 4.4% 0.180 10.83 372.89
-H-Gmb
-PPL 100.0% 0.349 100.29 54.04 100.0% 0.189 19.07 321.16 100.0% 0.163 10.33 384.04
-KNN-L2 8 52.1% 0.344 98.98 59.66 99.9% 0.222 25.00 280.97 23.6% 0.185 10.72 376.35
-KNN-CE 1 100.0% 0.368 86.49 68.27 100.0% 0.226 14.27 342.64 100.0% 0.173 3.17 435.04
-S-Gmb
-PPL 55.0% 0.826 173.59 9.61 100.0% 0.183 10.28 377.79 48.4% 0.386 45.61 167.14
-KNN-L2 8 96.3% 0.358 74.00 85.36 100.0% 0.201 10.17 373.83 100.0% 0.193 5.73 404.18
-KNN-CE 1 100.0% 0.569 194.74 16.60 100.0% 0.233 15.41 329.78 100.0% 0.186 3.83 425.33
-Softmax
-PPL 100.0% 1.112 309.60 4.55 81.6% 0.701 37.99 196.72 99.8% 0.725 83.07 77.98
-KNN-L2 1 100.0% 0.358 76.29 81.61 100.0% 0.224 12.08 354.36 100.0% 0.387 32.20 210.61
-2 100.0% 0.366 81.69 71.75 100.0% 0.226 12.04 348.48 100.0% 0.175 3.62 427.42
-4 100.0% 0.379 87.05 62.37 100.0% 0.205 9.01 383.56 100.0% 0.196 5.64 403.06
-8 99.9% 0.343 73.72 88.53 100.0% 0.199 12.64 361.23 100.0% 0.204 7.91 381.65
-KNN-CE 1 100.0% 0.366 79.80 74.69 100.0% 0.225 14.08 345.56 100.0% 0.175 2.81 437.72
-2 100.0% 0.380 106.05 50.75 100.0% 0.227 14.34 338.37 100.0% 0.187 4.16 418.57
-4 100.0% 0.748 255.59 9.03 100.0% 0.228 16.73 326.78 100.0% 0.204 6.98 387.88
-8 100.0% 1.240 467.86 3.01 100.0% 0.219 14.24 343.23 100.0% 0.199 5.27 404.10
-Table 1 reports the codebook usage and reconstruction quality scores—including root mean squared error (rMSE), Inception Score (IS; Salimans et al.,, 2016), and Fréchet Inception Distance (FID; Heusel et al.,, 2017)–for the proposed method (“KNN-L2/CE”) and the baselines.
-Reconstruction through softmax-smoothed quantization (i.e., without Gumbel randomness) deteriorated substantially when combined with the perplexity-based regularization (“PPL”; Equation 11).
-This degradation stems from the mismatch between soft quantization during the training and hard quantization at inference; the smoothed quantizers deviated from the simplex vertices, as reflected in their high individual perplexity, $\exp(-\sum_{m=1}^{M}p_{m}\log p_{m})$ (Table 2).
-These observations support the argument made in the Introduction that perplexity-based regularization alone does not promote tight smoothing.
+$\times16$$\times$32; 1024 $\times64$$\times$3; 8196 $\times64$$\times$32; 8196 $\times64$$\times$2048; 8196
+Method $K/4$ Code Use ($\uparrow$) rMSE ($\downarrow$) FID ($\downarrow$) IS ($\uparrow$) Code Use rMSE FID IS Code Use rMSE FID IS Code Use rMSE FID IS
+STE Euclid 4.5% 0.404 124.86 36.61 100.0% 0.167 7.25 402.84 1.7% 0.235 22.02 290.24 0.1% 0.274 32.60 227.73
+Cosine 3.0% 0.381 117.95 41.48 70.9% 0.197 13.97 348.31 2.8% 0.186 12.11 363.86 0.2% 0.227 23.03 290.15
+SimVQ 100.0% 0.340 87.26 71.33 100.0% 0.170 7.44 400.41 100.0% 0.148 3.97 436.29 70.8% 0.682 286.18 3.19
+RE Euclid 3.1% 0.460 170.30 19.25 78.85% 0.171 10.21 377.56 0.5% 0.271 40.58 203.52 0.1% 0.301 53.81 153.77
+Cosine 2.8% 0.423 157.77 24.33 99.5% 0.194 14.67 344.17 4.4% 0.180 10.83 372.89 1.3% 0.193 16.98 331.17
+HG PPL 100.0% 0.349 100.29 54.04 100.0% 0.189 19.07 321.16 100.0% 0.163 10.33 384.04 100.0% 0.165 8.89 392.77
+KNN-CE 1 100.0% 0.368 86.49 68.27 100.0% 0.226 14.27 342.64 100.0% 0.173 3.17 435.04 100.0% 0.177 3.42 432.08
+SG PPL 55.0% 0.826 173.59 9.61 100.0% 0.183 10.28 377.79 48.4% 0.386 45.61 167.14 34.6% 1.253 368.71 1.58
+KNN-CE 1 100.0% 0.569 194.74 16.60 100.0% 0.233 15.41 329.78 100.0% 0.186 3.83 425.33 100.0% 0.188 3.73 422.84
+Softmax PPL 100.0% 1.112 309.60 4.55 81.6% 0.701 37.99 196.72 99.8% 0.725 83.07 77.98 89.3% 1.345 84.52 80.39
+KNN-L2 1 100.0% 0.358 76.29 81.61 100.0% 0.224 12.08 354.36 100.0% 0.387 32.20 210.61 75.9% 1.259 104.22 56.87
+2 100.0% 0.366 81.69 71.75 100.0% 0.226 12.04 348.48 100.0% 0.175 3.62 427.42 58.3% 1.302 94.82 62.82
+4 100.0% 0.379 87.05 62.37 100.0% 0.205 9.01 383.56 100.0% 0.196 5.64 403.06 68.3% 1.283 96.45 73.56
+8 99.9% 0.343 73.72 88.53 100.0% 0.199 12.64 361.23 100.0% 0.204 7.91 381.65 80.8% 1.134 73.79 92.46
+KNN-CE 1 100.0% 0.366 79.80 74.69 100.0% 0.225 14.08 345.56 100.0% 0.175 2.81 437.72 100.0% 0.424 18.58 289.42
+2 100.0% 0.380 106.05 50.75 100.0% 0.227 14.34 338.37 100.0% 0.187 4.16 418.57 100.0% 0.212 4.22 416.56
+4 100.0% 0.748 255.59 9.03 100.0% 0.228 16.73 326.78 100.0% 0.204 6.98 387.88 100.0% 0.190 4.70 411.70
+8 100.0% 1.240 467.86 3.01 100.0% 0.219 14.24 343.23 100.0% 0.199 5.27 404.10 100.0% 0.201 3.77 416.71
+Table 1 reports the codebook utilization and reconstruction performance—including root mean squared error (rMSE), Inception Score (IS; Salimans et al.,, 2016), and Fréchet Inception Distance (FID; Heusel et al.,, 2017)–for the proposed method (“KNN-L2/CE”) and the baselines.
+Reconstruction quality deteriorated substantially when softmax-smoothed quantization (i.e., without Gumbel randomness) was combined with the perplexity-based regularization (“PPL”; Equation 11).
+This degradation arises from a mismatch between soft quantization during the training and hard quantization at inference; the smoothed quantizers deviated from the simplex vertices, as reflected by their high individual perplexity, $\exp(-\sum_{m=1}^{M}p_{m}\log p_{m})$ (Table 2).
+These results support the argument made in the Introduction that perplexity-based regularization alone does not promote tight smoothing.
 Table 2: Tightness of softmax-based smoothing (without Gumbel sampling), measured by the individual perplexity of smoothed quantizers, $\exp(-\sum_{m=1}^{M}p_{m}\log p_{m})$.
@@
 Feature Map Size; Codebook Size
-$\times16$$\times$32; 1024 $\times64$$\times$3; 8192 $\times64$$\times$32; 8192
-Method $K/4$ 75% 90% 99% Max 75% 90% 99% Max 75% 90% 99% Max
-PPL 910.40 933.12 963.29 995.91 7559.89 7561.99 7563.10 7565.02 7230.60 7267.06 7315.62 7424.16
-KNN-L2 1 1.17 1.59 2.36 6.26 1.00 1.00 1.00 4.00 2.24 40.45 86.53 249.64
-2 1.00 1.08 1.86 27.37 1.00 1.00 1.06 4.65 1.14 1.55 2.33 7.92
-4 1.00 1.06 1.86 62.99 1.00 1.00 1.00 4.61 1.00 1.06 1.80 20.42
-8 1.00 1.10 1.90 120.68 3.33 3.87 5.22 12.03 1.00 1.05 1.79 107.70
-KNN-CE 1 1.18 1.60 2.38 6.28 3.96 4.51 5.81 11.96 1.09 1.45 2.17 5.94
-2 1.23 1.72 2.73 8.07 4.04 4.66 6.16 11.77 1.06 1.38 2.08 6.15
-4 1.42 1.78 2.38 5.45 4.16 4.86 6.57 13.04 1.08 1.44 2.16 6.58
-8 2.38 2.80 3.71 7.78 4.46 5.31 7.39 13.81 1.29 1.75 2.65 9.26
-The perplexity-based regularization was only effective when combined with the Gumbel-softmax sampling, and onehot quantization in the forward computation by STE was necessary to ensure full codebook usage (“Hard-Gumbel”).
-Otherwise, code collapse was not prevented when the channel dimensionality was large ($C=32$; “Soft-Gumbel”).
-By contrast, the proposed KNN-based regularization successfully prevented code collapse and achieved near-complete codebook utilization without resorting to Gumbel-softmax sampling.
-The reconstruction quality was also superior or competitive with the Gumbel-softmax + perplexity approach across all feature map sizes.
+$\times16$$\times$32; 1024 $\times64$$\times$3; 8192 $\times64$$\times$32; 8192 $\times64$$\times$2048; 8192
+Method $K/4$ 75% 90% 99% Max 75% 90% 99% Max 75% 90% 99% Max 75% 90% 99% Max
+PPL 910.40 933.12 963.29 995.91 7559.89 7561.99 7563.10 7565.02 7230.60 7267.06 7315.62 7424.16 8191.71 8191.85 8192.02 8192.61
+KNN-L2 1 1.17 1.59 2.36 6.26 1.00 1.00 1.00 4.00 2.24 40.45 86.53 249.64 8191.66 8191.82 8191.98 8192.34
+2 1.00 1.08 1.86 27.37 1.00 1.00 1.06 4.65 1.14 1.55 2.33 7.92 8191.49 8191.72 8191.94 8192.35
+4 1.00 1.06 1.86 62.99 1.00 1.00 1.00 4.61 1.00 1.06 1.80 20.42 8191.63 8191.80 8191.97 8192.36
+8 1.00 1.10 1.90 120.68 3.33 3.87 5.22 12.03 1.00 1.05 1.79 107.70 8191.69 8191.82 8191.97 8192.41
+KNN-CE 1 1.18 1.60 2.38 6.28 3.96 4.51 5.81 11.96 1.09 1.45 2.17 5.94 1.90 124.24 503.28 1024.38
+2 1.23 1.72 2.73 8.07 4.04 4.66 6.16 11.77 1.06 1.38 2.08 6.15 1.28 2.24 16.83 676.31
+4 1.42 1.78 2.38 5.45 4.16 4.86 6.57 13.04 1.08 1.44 2.16 6.58 1.36 1.98 3.16 11.94
+8 2.38 2.80 3.71 7.78 4.46 5.31 7.39 13.81 1.29 1.75 2.65 9.26 1.86 2.51 4.28 26.69
+Figure 2: (Top) Histogram of per-image codebook utilization on the validation dataset, defined as the percentage of codes used at least once within each image.
+(Bottom) Histogram of code popularity, measured as the proportion of validation images that contain each code.
+Perplexity-based regularization was effective only when combined with the Gumbel-softmax sampling, and onehot quantization in the forward pass via STE was necessary to ensure full codebook utilization (“Hard-Gumbel”).
+Without hard quantization, code collapse was not prevented at higher channel dimensionalities ($C\geq 32$; “Soft-Gumbel”).
+By contrast, the proposed KNN-based regularization consistently prevented code collapse and achieved near-complete codebook utilization without resorting to Gumbel-softmax sampling.
+Reconstruction performance was also superior to, or competitive with, the Gumbel-softmax plus perplexity baseline across all feature map sizes.
 The choice of deviation metric (L2 distance vs.
-cross-entropy, CE) did not result in consistent global superiority, although noticeable code collapse occurred when KNN-L2 regularization was combined with hard Gumbel-softmax sampling, which nonetheless had limited effect on the reconstruction quality.
-In terms of computational efficiency, however, cross-entropy proved more favorable.
-Notably, its performance was highest when the number of neighbors was minimal ($K/4=1$) across all feature map sizes (indicated by the underlined scores in Table 1); increasing $K$ did not yield further improvements.
-Although most smoothed quantizers remained unregularized under this setting, they still achieved low individual perplexities after the training (Table 2), indicating a tight approximation of quantization.
-By contrast, minimizing only the L2 distance between the simplex vertices and their single nearest neighbor per GPU led to performance degradation for the feature map of size $64\times 64\times 8192$.
-Since requiring more neighbors necessitates a larger batch size (see §5.3 for details), cross-entropy emerges as a more scalable option.
-Finally, both STE and rotational gradient estimation (RE) exhibited severe code collapse when the channel dimensionality was large ($C=32$), reaffirming prior findings that reducing channel dimensionality is necessary for stable training (Yu et al., 2022a, ; Yu et al.,, 2024; Mentzer et al.,, 2024).
-By contrast, SimVQ maintained full codebook usage across all settings and achieved strong reconstruction quality even at high dimensionality.
-However, the next section presents a case study in which all gradient-estimation approaches—including SimVQ—encounter code collapse, highlighting the difficulty of achieving robust prevention of code collapse across different settings.
+cross-entropy, CE) did not exhibit a consistent global advantage.
+One notable exception occurred at extremely high channel dimensionality ($C=2048$).
+In this regime, KNN-L2 failed to achieve full codebook utilization, whereas KNN-CE maintained complete utilization.
+Moreover, the soft quantizers under KNN-L2 failed to concentrate around the simplex vertices in the high-dimensionality condition, as indicated by their near-maximal individual perplexity values (Table 2).
+This looseness in smoothing also led to degraded reconstruction quality (Table 1).
+Cross-entropy was additionally favorable in terms of computational efficiency.
+Its performance was highest when using the minimum number of neighbors ($K/4=1$) for nearly all feature map sizes (underlined entries in Table 1); increasing $K$ yielded no further gains except in the extreme case of $C=2048$.
+Although most soft quantizers remained formally unregularized under this setting, they nonetheless achieved low individual perplexity after training (Table 2), indicating a tight approximation to hard quantization.
+By contrast, minimizing the L2 distance to only a single nearest neighbor per GPU led to performance degradation for feature maps of size $64\times 64\times\{32,2048\}$.
+Because increasing the number of neighbors requires a larger batch size (see §5.3 for details), cross-entropy emerges as the more scalable option.
+Finally, both STE and rotational gradient estimation (RE) exhibited severe code collapse at higher channel dimensionalities ($C\geq 32$), reaffirming prior findings that reducing channel dimensionality is critical for stable training (Yu et al., 2022a, ; Yu et al.,, 2024; Mentzer et al.,, 2024).
+SimVQ was more robust, maintaining full codebook utilization and strong reconstruction performance at $C=32$.
+However, when the channel dimensionality increased to $C=2048$, its performance fell below that of classical STE-based models.
+Although SimVQ still utilized 70.8% of the codebook overall, each individual image was encoded by using no more than 1% of the codes (Figure 2).
+Effective vector quantization instead requires diverse code usage within individual images (typically 10–40%), a property that KNN-CE regularization preserves across channel dimensionalities ($C\in\{3,32,2048\}$).
+The next section presents a case study in which all gradient-estimation approaches—including SimVQ—exhibit more pronounced code collapse, highlighting the difficulty of achieving robust prevention of code collapse across different settings.
 4.2 Contrastive Learning
@@
 Table 3: Codebook usage in Wav2Vec 2.0 pretraining, evaluated on the LibriSpeech dev-clean split (similar results were observed for the other dev/test splits).
-The proposed method is denoted as “KNN-L2/CE”.
 #Codebooks $\times$ Codebook Size
@@
 Method $K/4$ Codebook#1 Codebook#2
-STE
-Euclid 0.8% 0.9% 0.9%
+STE Euclid 0.8% 0.6% 0.6%
 Cosine 0.2% 0.6% 0.6%
 SimVQ 0.2% 0.6% 0.6%
-RE
-Euclid 2.5% 0.6% 0.6%
+RE Euclid 2.5% 0.6% 0.6%
 Cosine 0.2% 0.6% 0.6%
-H-Gmb
-PPL 0.7% 0.6% 0.6%
-KNN-L2 2 0.2% 0.6% 0.6%
-KNN-CE 2 99.7% 100.0% 100.0%
-S-Gmb
-PPL 0.3% 0.6% 0.6%
-KNN-L2 2 90.1% 0.6% 0.6%
-KNN-CE 2 100.0% 100.0% 100.0%
-Softmax
-PPL 0.2% 1.2% 1.2%
-KNN-L2 1 82.4% 100.0% 100.0%
-2 60.4% 100.0% 100.0%
-KNN-CE 1 99.5% 100.0% 100.0%
-2 100.0% 100.0% 100.0%
+HG PPL 0.7% 0.6% 1.2%
+KNN-L2 2 0.2% 0.6% 0.9%
+KNN-CE 2 99.7% 70.0% 100.0%
+SG PPL 0.3% 0.6% 0.9%
+KNN-L2 2 90.1% 21.6% 99.1%
+KNN-CE 2 100.0% 54.1% 100.0%
+Softmax PPL 0.2% 0.6% 0.6%
+KNN-L2 1 82.4% 89.7% 100.0%
+2 60.4% 87.2% 100.0%
+KNN-CE 1 99.5% 76.2% 100.0%
+2 100.0% 76.6% 100.0%
 Models were trained on the LibriSpeech dataset, combining all training splits (train-clean-100 + train-clean-360 + train-other-500; Panayotov et al.,, 2015).
@@
 Remarkably, SimVQ—despite achieving full codebook usage in the discrete autoencoding experiments—offered no observable benefit in this learning paradigm.
-By contrast, the proposed KNN-based regularization ensured (near-)complete code utilization in both conditions when cross-entropy was used as the divergence metric on the simplex.
+By contrast, the proposed KNN-based regularization ensured high code utilization in both conditions.
 As in the discrete autoencoding experiments, a single neighbor per GPU was sufficient to obtain this effect.
-The L2 metric, on the other hand, proved insufficient to prevent code collapse, especially when combined with Gumbel sampling.
 5 Discussions
@@
 Acknowledgments
-Hidden for double-blind review.
+This study was supported by JSPS Grant-in-Aid for Scientific Research A (JP24H00774), B (JP22H03914), and C (JP24K15087); JST AIP Accelerated Program (JPMJCR25U6) and Core Research for Evolutional Science and Technology (JPMJCR22P5); and Kayamori Foundation of Informational Science Advancement (K35XXVIII620).
+The author also gratefully acknowledges the support of the Academic Center for Computing and Media Studies, Kyoto University, regarding the use of their supercomputer system.
 Appendix A Implementation of the Quantization Methods
@@
 Table 4: Hyperparameters for discrete autoencoding.
-Feature Map Size $16\times 16\times 32$ $\times 64\times 3 64×64$$\times 64\times 32$
-Codebook Size 1024 8196 8196
-Latent Channels $128\to 128\to 64\to 64\to 32\to 32$ $128\to 64\to 32\to 3$ $128\to 64\to 32\to 32$
-Height & Width $256\to 128\to 64\to 32\to 16\to 16$ $256\to 128\to 64\to 64$ $256\to 128\to 64\to 64$
-Batch Size 64 64 64
-Training Epochs 25 20 20
-Warmup Iterations 16,000 16,000 16,000
+Feature Map Size $16\times 16\times 32$ $\times 64\times 3 64×64$$\times 64\times 32$ $64\times 64\times 2048$
+Codebook Size 1024 8196 8196 8196
+Latent Channels $128\to 128\to 64\to 64\to 32\to 32$ $128\to 64\to 32\to\{3,32,2048\}$
+Height & Width $256\to 128\to 64\to 32\to 16\to 16$ $256\to 128\to 64\to 64$
+Batch Size 64 64 64 64
+Training Epochs 25 25 20 20
+Warmup Iterations 16,000 16,000 16,000 16,000
 Inception Score (IS; Salimans et al.,, 2016) and Fréchet Inception Distance (FID; Heusel et al.,, 2017) were estimated using the ImageNet-pretrained Inception V3 provided in torchvision.
@@
 Both the convolutional encoder and Transformer were implemented using the publicly available code in torchaudio, and only the quantization module was implemented from scratch.
-All components are provided in the supplementary material.
 Input waveforms were randomly cropped to the length of 250k samples.